National Electric Company (NEC) is considering a $45.13 million project in its power systems division. Tom Edison, the company’s chief financial officer, has evaluated the project and determined that the project’s unlevered cash flows will be $3.24 million per year in perpetuity. Mr. Edison has devised two possibilities for raising the initial investment: Issuing 10-year bonds or issuing common stock. The company’s pretax cost of debt is 7.3 percent and its cost of equity is 12.1 percent. The company’s target debt-to-value ratio is 85 percent. The project has the same risk as the company’s existing businesses and it will support the same amount of debt. The tax rate is 23 percent.

Calculate the weighted average cost of capital. (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

Calculate the net present value of the project.
Answer #1

Cost of debt = 7.3% * (1-23%) = 5.621%

Cost of equity = 12.1%

WACC = Weight of debt * cost of debt + Weight of equity * Cost of equity

= 85% * 5.621% + 15% *12.1% = 6.59285%

Present value of the project = $3,240,000/ 6.59285% = $49,144,148.59

Net Present value = 49,144,148.59 - 45,130,000 = $4,014,148.59